OpenStudy (anonymous):

Sam, age 40, wants to retire at age 60. He estimates he will need $1,500,000 to retire. How much money should he invest now, if he can find an account that pays 2.75% interest, compounded weekly?

OpenStudy (ash2326):

@erica1 This is a problem of Compound Interest, do you know the formula for it?

OpenStudy (anonymous):

i dont :(

OpenStudy (anonymous):

Hint: 20 years x 52 weeks means you will have 1040 compounding periods.

OpenStudy (anonymous):

Also, your interest rate for each week is 1/52 of the yearly interest rate. Are you following so far? If you are, what is your equation? Are you able to set it up now?

OpenStudy (ash2326):

OK its formula is given as \[A=P(1+\frac r n)^{tn}\] A=total amount after the end of the period= here it's $ 1,500, 000 P=initial amount r= rate in decimal = 2.75 /100=0.0275 t= no. of years= 60-40 =20 years n= no. of times the amount is compounded in a year= =52 Could you plugin these in the formula to find P?

OpenStudy (anonymous):

@ash2326 , you are off on r. You have to divide r by 52 (in addition to your 100) to get the weekly interest rate

OpenStudy (anonymous):

i have no idea how to figure it out. i got 20.54?

OpenStudy (anonymous):

Ok, ash, I see your division in the formula. You're ok.

OpenStudy (anonymous):

Just plug the #'s in. The formula is there for you already.

OpenStudy (anonymous):

ok i did and i told you what i got

OpenStudy (anonymous):

You know your arithmetic is off because there is no way you can invest $20.54 and get 1500000 after 20 years.

OpenStudy (ash2326):

@erica1 20.54 is very small value Let's plugin the values \[1500000=P(1+\frac{0.0275}{52})^{52\times 20}\] \[1500000=P(1+\frac{0.0275}{52})^{1040}\] \[1500000=P\times 1.733\] Find P from this

OpenStudy (anonymous):

its the percentage not dollar amount

OpenStudy (ash2326):

@erica1 but we need to find the amount

OpenStudy (anonymous):

i thought you could use that percentage of the amount to find your down payment or whatever

OpenStudy (anonymous):

P stands for Principal, not Percentage.

OpenStudy (anonymous):

i know

OpenStudy (ash2326):

so erica1 did you understand this?

OpenStudy (anonymous):

i got 25,995

OpenStudy (ash2326):

@erica1 check your calculations once more, this is also not correct

OpenStudy (anonymous):

i cant do it. never mind. thanks anyway guys

OpenStudy (ash2326):

@erica1 Don't give up, it's pretty easy. Check your calculations again and you'd get the answer

OpenStudy (anonymous):

i dont know how to check it. i dont even know how to do this stuff.

OpenStudy (anonymous):

i took 1,500,000 x 1.733%

OpenStudy (ash2326):

We have \[1, 500, 000=P\times 1.733\] no percent it's a division problem

OpenStudy (anonymous):

865,551.06

OpenStudy (ash2326):

Yeah, you're right But I can see that you didn't try to understand. Would you want to do that now?

OpenStudy (anonymous):

thank you

OpenStudy (ash2326):

Would you like to?

OpenStudy (anonymous):

i wont be able to remember it. im just trying to finish this test. it was due 2 weeks ago

OpenStudy (ash2326):

@erica1 you got the answer for this, but this won't help you. Don't seek just answers, try to understand the concept. Please don't ask for help here, while you are taking a test. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

thats what everyone does

OpenStudy (ash2326):

You don't have to follow everyone.

OpenStudy (anonymous):

this is all i need to graduate. but i cant do that if i dont pass. id like to have a high school diploma this december but i do online school and no one else in my family knows how to do this crap

OpenStudy (amistre64):

without knowing the criteria for you test taking, its hard to pass judgement. There are teachers which allow open book, open note tests; there are those that allow take home tests as well. There are a lot of teachers of the mindset that if you can work out a solution, then the means are not that important. The key is in gaining a better understanding of the material and avoiding the "just give an answer" mentality.